40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 13 21:30:55 2020, shruti, Update, ALS, ALS OOL noise with arms locked 20200213_ALS.pdf
    Reply  Fri Feb 14 14:02:13 2020, shruti, Update, ALS, ALS OOL noise with arms locked 819472_10.pdf225642_10.pdf693500_10.pdf
       Reply  Fri Feb 14 14:52:41 2020, gautam, Update, ALS, ALS OOL noise with arms locked 
          Reply  Tue Feb 18 18:14:59 2020, shruti, Update, ALS, ALS OOL noise with arms locked ALSNoise.pdf
             Reply  Wed Feb 19 22:20:22 2020, rana, Update, ALS, ALS OOL noise with arms locked 
             Reply  Fri Feb 21 10:59:08 2020, shruti, Update, ALS, PDH error signals? IMG_0082.jpgIMG_0083.jpgIMG_0084.jpgIMG_0085.jpg
                Reply  Fri Feb 21 13:02:53 2020, Koji, Update, ALS, PDH error signals? 
                   Reply  Fri Feb 21 20:44:18 2020, shruti, Update, ALS, ALS OOL noise and PDH ALSNoise.pdfIMG_0086.jpg
                      Reply  Sun Feb 23 18:15:22 2020, rana, Update, ALS, ALS OOL noise and PDH 
Message ID: 15216     Entry time: Tue Feb 18 18:14:59 2020     In reply to: 15214     Reply to this: 15217   15218
Author: shruti 
Type: Update 
Category: ALS 
Subject: ALS OOL noise with arms locked 

We proceeded with the trying to measure the ALS out-of-loop noise of the X arm when the X arm cavity is green-locked using different PDH sideband frequencies.

Before doing the experiment, Koji helped us with getting the arm cavities locked in IR using LSC (length) and ASC (angular).

With the arms locked in IR and green, we repeated the same measurements as before at different sideband frequencies (Refer Attachment 1 - label in Hz). We did not optimize the phase nor did we look at the PDH error signal today which is possibvly why we did not see an improvement in the noise. We will look into this possibly tomorrow.

Attachment 1: ALSNoise.pdf  569 kB  Uploaded Tue Feb 18 20:20:27 2020  | Hide | Hide all
ALSNoise.pdf
ELOG V3.1.3-